# Guided Diffusion for Fast Inverse Design of Density-based Mechanical Metamaterials

### v1

@@ Results
During sampling, we start with Gaussian noise and progressively denoise to obtain a clear structure. DDIM reverse sampling, on the other hand, can deduce the corresponding noise from a clear structure. From this perspective, noise can be considered as a latent space for microstructures. We perform spherical interpolation on the noise to obtain a series of continuously varying microstructures. Specifically, given the start and end structure for interpolation, we first determine their corresponding Gaussian noises, denoted as $x_{1}$ and $x_{2}$. Subsequently, we generate a series of Gaussian noises $cos(\theta)x_{0}+sin(\theta)x_{1}$ where $\theta$ sweeps from 00 to $\frac{\pi}{2}$. By using this series of noises for unconditional generation, a set of microstructures with continuous variations in both shape and physical properties can be obtained. Figure 5 illustrates two examples of interpolation. In the first example, we chose microstructures from the dataset with extreme Young’s modulus as the start and end points. The interpolation trajectory is shown in Figure 5c, while Figure 5a displays interpolated microstructures and their modulus spheres, demonstrating a smooth transition in both shape and physical properties.
However, similar to the results in previous articles [41], the interpolation trajectory obtained solely through noise interpolation is uncontrollable. The resulting microstructures cannot be directly applied to multi-scale topology optimization. Therefore, we design a family of microstructures whose bulk modulus is close to HS bound. Figure 5d illustrates the interpolation trajectory of this microstructure family, 11 selected microstructures and their modulus spheres are depicted in Figure 5b. We trained a regressor to predict the ratio of Bulk modulus to the HS bound. We then used this regressor with classifier-guidance conditions for generation, replacing the unconditional generation in the original interpolation method. This ensures that our interpolated microstructures all possess extreme Bulk modulus. Additionally, we found it challenging to obtain a dense microstructure family through a single interpolation. Even with denser interpolation points, we only get more similar microstructures, and there are always some sections cannot be filled. As a remedial measure, we initially interpolated 64 microstructures. Subsequently, we continued the interpolation using the ends of the gaps as start and end points, resulting in the interpolation of 16 additional microstructures. This iterative process was repeated until it yielded a dense microstructure family. The final microstructure family we obtained covers almost all volume fractions between the start and end points of the initial interpolation, and all microstructures possess extreme Bulk modulus.
Figure 6: Mechanical cloak via multi-scale design. Four mechanical invisibility cloaks. The left two consider cutting out a cylinder in a cube and designing an invisibility cloak around it. The right two cut out a rectangular column in the cube and design an invisibility cloak around it. Middle sections and local $3\times 3\times 3$ structures are rendered. The cyan structures are the reference structures.
@@
To exemplify our proposed method, we employ voxel-based datasets to design mechanical cloak multiscale systems applications. The reference structures (absent voids and cloak) consist of $40\times 40\times 40$ periodically tessellated base cells selected from the metamaterials dataset. We design the invisibility cloak for two situations: (1) The first case involves excising a cylinder with a radius of $\frac{20}{3}$, and subsequently filling the outer circular region (comprising a cylinder with an outer radius of $\frac{40}{3}$) with material to achieve the design of the invisibility cloak. (2) the second case, the cylinder from the first case is substituted with a rectangular column. The cross-sectional shape of this column is a square with a side length equal to the diameter of the original cylinder. Simultaneously, we explored applying two distinct reference structures: one characterized by a plate structure with a thickness of eight voxels on each face and the other representing an isotropic structure identified from the dataset to emulate materials with isotropic properties. The effectiveness of the cloak’s stealth capabilities is considered superior when the discrepancy between the displacement and the reference displacement is minimized. In all four instances, a noteworthy reduction in displacement distortion was achieved.
@@ Discussion
This paper introduces a novel and efficient solution through a fast inverse design method leveraging an advanced deep generative AI algorithm. The core of this approach is the self-conditioned diffusion model, a powerful tool capable of generating voxel-based mechanical metamaterials with a resolution as high as $128^{3}$. What sets our method apart is its remarkable ability to approach the specified homogenized tensor matrix in 3 seconds. The immediate application of this rapid reverse design tool is evident in its capability to facilitate the exploration of extreme metamaterials. By providing a quick and efficient means to generate metamaterial structures, our approach allows for sequence interpolation in metamaterials, opening avenues for designing materials with tailored mechanical responses. Furthermore, the tool enables the generation of diverse microstructures, facilitating multi-scale design considerations.
In the realm of structural engineering and other mechanical systems, the flexibility and adaptability of our generative tool hold immense value. Beyond the immediate applications discussed, the tool is expected to stimulate further research by providing researchers with a swift and efficient means of exploring and testing novel mechanical metamaterial designs. In conclusion, our proposed fast inverse design method, driven by an advanced deep generative AI algorithm, represents a significant advancement in the field of mechanical metamaterials. Its capabilities offer not only immediate practical applications but also the potential to inspire and drive future research endeavors in materials science and engineering.
@@ Methods
All density-based microstructures are obtained by the inverse homogenization method in a cube domain. The cube is discretized into $n$ voxel elements, with each element assigned a density value of 00 (void) or $1$ (solid). Various combinations of 00 and $1$ give rise to different microstructures, totaling $2^{n}$ combinations. The number of microstructure variations expands exponentially with the resolution parameter, $n$. Consequently, voxel-based representations exhibit significant representational capabilities. However, it is important to note that not all combinations yield satisfactory microstructures, and the huge amount of data may impose a considerable burden on subsequent training and machine learning algorithms. When constructing the data set, we take into account the cube symmetry employing the $\frac{1}{8}$ unit cell lattice to simplify the representation. Subsequently, symmetric operations are applied to the $\frac{1}{8}$ unit cell lattice. This process ultimately leads to a cubic symmetric microstructure to facilitate calculation. To create a sufficiently large design space for a metamaterial database based on voxels, we establish microstructures at a resolution of $128^{3}$ and subsequently store $\frac{1}{8}$-th of it, equivalent to $64^{3}$ microstructures.
Extension from perturbation data
Our perturbation algorithm is an iterative process wherein, at each step, we locally apply elastic deformation to specified structures. The deformation field is characterized by random Gaussian noise. Initially, this perturbation is broadly applied to the entire dataset. As the iterations progress, each subsequent step specifically targets a range of Poisson’s ratio that is underrepresented in the current dataset. This targeted approach enables us to concentrate the perturbation on areas where modulus coverage is less extensive, thereby enhancing the dataset’s diversity in these particular aspects. Through continuously adjusting our focus to different Poisson’s ratio ranges with each iteration, we gradually achieve a dataset with a broad and comprehensive range of modulus coverage.
@@
The objective functions for optimization are the bulk modulus, shear modulus, and Poisson’s ratio. We impose volume constraints during the optimization process. The total optimization results consist of 10,000 structures for volume modulus and shear modulus, while 5,000 structures are obtained for Poisson’s ratio. The volume fractions considered in the optimization range from 0.2 to 0.9. In utilizing the iterative perturbation method, a total of 60,000 structures were initially generated. Following the removal of disconnected portions within the microstructure, as illustrated in Figure 1, 74,717 microstructures remained. To enhance training efficiency, a subset of 10,000 data was uniformly selected from the original dataset comprising 74,717 entries.
Framework of diffusion model
The backbone denoising network is an U-Net based on the standard 3D convolutional neural network. Following LAS-Diffusion [46], the U-Net is composed of 5 levels: $64^{3},32^{3},16^{3},8^{3}$, and $4^{3}$, with feature dimensions gradually increasing to 32, 64, 128, 256, and 256, respectively. Each level is constructed with a ResNet block, which includes two convolution layers with a kernel size of 3. In the U-Net’s bottleneck, two ResNet blocks are introduced. To map the voxel features at the finest level to a surface-occupancy value, a convolution layer is added at the conclusion of the network.

### v2

@@ Results
During sampling, we start with Gaussian noise and progressively denoise to obtain a clear structure. By reversing this process, we can deduce the corresponding noise from a clear structure. From this perspective, noise can be considered a latent space for microstructures. We perform spherical interpolation on the noise to obtain a series of continuously varying microstructures. Specifically, given the start and end structure for interpolation, we first determine their corresponding Gaussian noises, denoted as $x_{1}$ and $x_{2}$. Subsequently, we generate a series of Gaussian noises $cos(\theta)x_{0}+sin(\theta)x_{1}$ where $\theta$ sweeps from 00 to $\frac{\pi}{2}$. By using this series of noises for unconditional generation, we obtain a set of microstructures that smoothly change in both geometry and physical properties.
Figure 5a shows an interpolation application case with start and end having extreme Young’s moduli, and the corresponding elastic surfaces are shown in Figure 5d. For such an extreme start and end, a smoothly transitioned sequence can be obtained, demonstrating the robustness of the interpolation method. Although we have not introduced physical information into the unconditional generative model, the noise of smooth transition leads to geometric shapes of smooth transition, resulting in their modulus also being smoothly transitional. However, such an interpolation solely for noise is not controllable for interpolation trajectories in the property space. This is consistent with past results [41]. As a result, there has always been a gap between microstructure interpolation and multiscale design.
As a further effort, we slightly modify the interpolation algorithm and obtained a family of microstructures with bulk moduli close to the HS bound and volume fractions ranging from 0.07 to 0.8, as shown in Figure 5c and Figure 5e. We first obtain a series of microstructures with relatively high bulk moduli by selecting the start and end microstructures with bulk modulus values close to the HS bound, as shown in Figure 5b. However, the bulk modulus of several microstructures falls below 86% of the HS bound. Then, following the so-called universal guidance [50], we add two controls for high bulk modulus and boundary similarity during the sequence generation and obtained the sequence shown in Figure 5c. Its bulk modulus and transition smoothness are superior to those in Figure 5b.
Figure 6: Mechanical cloak via multiscale design. a Problem definition. b C11 distribution of optimized result. c A cross-section of the optimized result. d The range of elasticity tensors that can be generated by the generative model. e A 3*3 part of the optimized result.
@@
We also employ our generative model for multiscale design. By trivially extending the optimization problem and methods for 2D mechanical cloaks [51] to 3D, we design mechanical cloaks in 3D. The reference structures (absent voids and cloaks) consist of $40\times 40\times 40$ periodically tessellated base cells selected from the metamaterial dataset. We design an invisibility cloak for such a situation by excising a cylinder with a radius of $\frac{20}{3}$ and subsequently filling the outer circular region (comprising a cylinder with an outer radius of $\frac{40}{3}$) with material to design the invisibility cloak. Our surrounding structure is an isotropic structure identified from the dataset to emulate materials with isotropic properties, such as soil. Our designed invisibility cloak can reduce displacement distortion from 34.3% to 4.12%, achieving effective cloaking performance.
@@ Discussion
This paper introduces a fast inverse design method via an advanced deep generative AI algorithm to generate metamaterials matching a specified tensor. A high-resolution voxel-based dataset is generated through optimization and active learning methods. Its broad coverage facilitates a comprehensive exploration of mechanical metamaterials. The self-conditioned diffusion model, the heart of our approach, is a powerful tool that can rapidly generate voxel-based mechanical metamaterials with a resolution as high as $128^{3}$. Its exceptional ability to approach the specified homogenized tensor matrix in just 0.42 seconds distinguishes our method. This rapid inverse design tool has immediate applications in exploring extreme metamaterials, enabling quick and efficient generation of metamaterial structures. This approach also opens up possibilities for sequence interpolation in metamaterials, paving the way for designing materials with tailored mechanical responses. Additionally, the tool facilitates the generation of diverse microstructures, enhancing multiscale design considerations.
Our generative tool, with its rapid design capabilities, holds immense value not only in structural engineering but also in other mechanical systems. Reducing the number of optimization cycles decreases the time and resource consumption, enhances the research efficiency, and renders metamaterial development more cost-effective. Its potential impact extends to the biomedical domain, where it can significantly shorten the customization cycle for bone implants, benefiting patients. In conclusion, our proposed fast inverse design method, driven by an advanced deep generative AI algorithm, significantly advances the field of mechanical metamaterials and has the potential to inspire and drive future research endeavors in a wide range of fields, including materials science and engineering.
@@ Methods
Density-based microstructures are obtained by the inverse homogenization method in a cube domain. The cube is discretized into $n$ voxel elements, with each element assigned a density value of 00 (void) or $1$ (solid). Various combinations of 00 and $1$ give rise to different microstructures, resulting in $2^{n}$ combinations. The number of microstructure variations expands exponentially with the resolution parameter $n$. Consequently, voxel-based representations exhibit significant representational capabilities. However, it is important to note that not all combinations yield satisfactory microstructures, and the large amount of data may impose a considerable burden on subsequent training and machine learning algorithms. When constructing the dataset, we consider the cube symmetry employing $\frac{1}{8}$ of unit cell lattice to simplify the representation. Subsequently, symmetric operations are applied to the $\frac{1}{8}$ unit cell lattice. This process ultimately leads to a cubic symmetric microstructure to facilitate calculation. To create a sufficiently large design space for a metamaterial database based on voxels, we establish microstructures at a resolution of $128^{3}$ and subsequently store $\frac{1}{8}$-th of it, equivalent to $64^{3}$ microstructures.
@@
The LIVE3D framework [22] is used to generate the initial dataset. The objective functions are the bulk modulus, shear modulus, and Poisson’s ratio with volume constraints. The total optimization results consist of 5,000 structures each for bulk modulus and shear modulus, while 4,396 structures are obtained targeting Poisson’s ratio. The volume fractions considered in the optimization range from 0.05 to 1. After iterative augmentation via active learning, our final dataset comprises 144,054 microstructures, as illustrated in Figure 1. However, the distribution of these 144,054 data points in the C11-C12-C44-vol space is uneven. To eliminate bias and improve training efficiency, we remove some data from overly dense regions, resulting in a training set containing 21,212 data points.
Framework of the diffusion model
The backbone denoising network is a U-Net based on the standard 3D convolutional neural network. Following LAS-Diffusion [52], U-Net is composed of 5 levels: $64^{3},32^{3},16^{3},8^{3}$, and $4^{3}$, with feature dimensions gradually increasing to 32, 64, 128, 256, and 256, respectively. Each level is constructed with a ResNet block, which includes two convolution layers with a kernel size of 3. In the U-Net bottleneck, two ResNet blocks are introduced. To map the voxel features at the finest level to a surface occupancy value, a convolution layer is added at the conclusion of the network.
Universal guidance for interpolation
Specifically, universal guidance suggests that the classifier in classifier guidance can be replaced with any objective function, and its gradient can be used to guide the generation. The objective function we use for Figure 5c is:
$E(\hat{x}_{0,i})=-3f(p(\hat{x}_{0,i}))+E_{boundary}(p(\hat{x}_{0,i}),x_{0,i-1})+\alpha E_{boundary}(p(\hat{x}_{0,i}),x_{0,start})+(1-\alpha)E_{boundary}(p(\hat{x}_{0,i}),x_{0,end}).$
Here, $\hat{x}_{0,i}$ is the predicted clear microstructure in the reverse denoising process for the ith microstructure in the interpolation sequence. $p(x)=\frac{tanh(64x+1)}{2}$ is a projection function to force $x$ to a 0-1 tensor. $f$ is a regressor that predicts the ratio of the bulk modulus of the microstructure to its upper limit at the volume fraction. The regression-derived $R^{2}$ is 0.99. $E_{boundary}(x,y)=||x[0,:,:]-y[0,:,:]||_{2}$, which measures the boundary difference between two microstructures $x$ and $y$. The last three terms of the objective function correspond to the boundary differences between the currently predicted microstructure in the denoising process and the previous microstructure in the interpolation sequence, the initial microstructure, and the final microstructure. $\alpha=\frac{i}{N}$, where $N$ is the length of the interpolation sequence.
Multiscale design
Our multiscale optimization method is an almost trivial extension of the 2D case [51]. A key difference is that they use microstructures from the dataset to fill the design space, whereas we use the generated microstructures. Therefore, similar to their use of property space constraints to restrict the design variables within the dataset range, we also need constraints to control the design variables within the generable range. We consider the generable range to be the coverage of the final version of the dataset. Specifically, for every element $e$ in the design space, there is a constraint: $\varphi(\mathbf{C}^{e})\geq 0$ to constrain the elasticity tensor $\mathbf{C}$ within the coverage of the final version of the dataset. This results in $40\times 40\times 40$ constraints. As suggested by Wang et al. [51], a single constraint is used to approximate these constraints. However, since our problem scale is much larger than that of their 2D problem, this approximation becomes less accurate, and the $40\times 40\times 40$ constraints are not always fully satisfied. Therefore, at the end of each optimization step, we project the elements that fall outside the dataset coverage to their best approximations within the dataset coverage. This ensures that all the elements can be accurately generated.
Data availability
The training data, which include voxel-based microstructures and their effective homogenized properties, as well as the generated data optimized to approach the extreme and the interpolated microstructure sequences generated in this study, have been deposited in Google Drive.
Code availability
The code used to train the generative modeling framework and obtain the inverse design of voxel-based microstructures has been uploaded to GitHub.